OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (−3, −4), B (−6, 2), C (0, 0), and D (−4, −2). A. 14 square units B. 15 square units C. 18 square units D. 20 square units

OpenStudy (anonymous):

Use the distance formula which is: \[\sqrt{(x _{1}-x _{2})^{2} \times (y _{1}-y _{2})^{2}}\]

OpenStudy (anonymous):

I'm not sure I understand....Where does the plot points come in? @irockout

OpenStudy (anonymous):

Ok so for example, if you were trying to find the base of the triangle (which is AB I believe) you would do: \[\sqrt{(-3--6)^{2}+ (-4-2)^{2}}\]

OpenStudy (anonymous):

So, for the problem i'm doing, CD, I'd do. \[\sqrt{(0 - -4)^{2} + (0 - -2)} \]

OpenStudy (anonymous):

yes except you also have to square the (0--2)

OpenStudy (anonymous):

so, \[\sqrt{(4^2 + 2^2}\]

OpenStudy (anonymous):

\[\sqrt{ 16 + 4}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{20} = 4 ??\]

OpenStudy (anonymous):

\[\sqrt{20} = 2\sqrt{5}\]

OpenStudy (anonymous):

4.47...

OpenStudy (anonymous):

But that's not one of the answers. @irockout

OpenStudy (anonymous):

are you sure your coordinates are correct?

OpenStudy (anonymous):

Yes. I'm positive.